# Do LLMs Have Political Correctness? Analyzing Ethical Biases and Jailbreak Vulnerabilities in AI Systems

### v1

@@ 4.2 Intensional Bias Analysis
Interestingly, the latest model, GPT-4o, exhibits higher jailbreak success rates, with a significant difference observed between marginalized and privileged groups. In particular, GPT-4o shows a difference of 0.128 between marginalized and privileged keywords, highlighting the substantial gap influenced by injected biases. Furthermore, most models demonstrate lower jailbreak success rates for prompts using privileged keywords compared to the baseline, whereas prompts using marginalized keywords exhibit higher success rates.
Intriguingly, the LLaMA3 model demonstrates a lower propensity for successful jailbreak attempts compared to other models. This lower rate can be attributed to Meta’s focus on creating more robust LLMs that are resistant to such vulnerabilities. It showcases Meta’s commitment to mitigating LLM safety risks through comprehensive alignment techniques (Vidgen et al., 2021; Ghosh et al., 2024; Caselli et al., 2020).
4.3 Performance Improvement using Bias-based Method
The bias-based approach we analyzed can also be applied to existing models. We confirmed performance improvement by applying the PCJailbreak method to Adaptive Attacks (Andriushchenko et al., 2024), the current state-of-the-art (SOTA) jailbreak attack. Specifically, for the llama2 model(Touvron et al., 2023), the performance improved from 98% to 100%, resulting in a 2% increase. For the phi model(Abdin et al., 2024), the performance enhanced from 95% to 99%, resulting in a 4% increase. These results show that the method can be easily integrated with existing techniques. The experiment was held by using the jailbreak attack artifact from JailbreakBench(Chao et al., 2024), which has 100 samples of Adaptive Attacks conversation.
@@ 4.3 Performance Improvement using Bias-based Method
Model Adaptive Attacks Adaptive Attacks with PCJailbreak
@@
4.4 Experimental Validation of PCDefense
To validate our defense method PCDefense, we conducted experiments on various models, including Llama2, Phi, and Qwen2. We observed the impact of our approach on Marginalized Rate Success and Privileged Rate Success metrics before and after applying our defense technique.
@@ 4.4 Experimental Validation of PCDefense
Table 4: Jailbreak Prevention performance of PCDefense
@@
4.5 Principle Component Analysis
In this section, we conduct a detailed analysis of how models interpret PCJailbreak prompts by performing Principal Component Analysis (PCA) (F.R.S., 1901) on the embeddings of Phi-mini, Qwen2, and LLaMA3 models, as shown in Appendix Figures 5, 6, 7 respectively. Our findings reveal that models with a high success rate in PCJailbreak tend to cluster PCJailbreak prompts together with benign prompts, while those with a lower success rate demonstrate the opposite behavior. The PCA results indicate that when using PCJailbreak, models tend to interpret it more similarly to benign prompts, with the degree of this similarity aligning with our observed PCJailbreak success rates.
@@ 5 Conclusion
In conclusion, our study highlights the complexities and potential unintended consequences of aligning large language models (LLMs) with safety measures aimed at preventing harmful outputs. The introduction of intentional biases for ethical behavior, while crucial for ensuring responsible AI, has led to significant discrepancies in jailbreak success rates based on gender and racial keywords. This discrepancy, coined as ”PCJailbreak,” underscores the risks that arise from safety-induced biases, particularly in terms of fairness and equality. Our findings emphasize the need for LLM developers to carefully balance safety and fairness in their models. Additionally, our proposed method of adding defense prompts without requiring additional inference or models shows a simple and scalable solution to mitigate jailbreak attempts. This simplicity and scalability of migation emphasizes the responsibility of LLM developers to adopt more comprehensive and proactive measures to address safety risks. Future work should focus on designing more inclusive and transparent alignment strategies to address the inherent challenges of AI safety while minimizing bias.
@@ Appendix A Appendix
Figure 5: PCA of Phi-mini model, which has 0.4386 PCJailbreak success rate. PCJailbreak samples are closely clustered with Benign samples.
@@
Figure 6: PCA of Qwen2 model, which has 0.1971 PCJailbreak success rate. PCJailbreak samples are relatively far from both harmful and benign samples.
@@
Figure 7: PCA of LLaMA3 model, which has 0.0650 PCJailbreak success rate. PCJailbreak samples are relatively close to harmful samples.

### v3

@@ 4.2 Ethical Bias Analysis
As shown in Table 2, GPT-4o has higher jailbreak success rates, with a notable 0.128 difference between marginalized and privileged groups. Most models show lower success rates for privileged keywords and higher rates for marginalized ones compared to the baseline, where the baseline refers to the prompt in Equation 1 without keywords. Intriguingly, the LLaMA3 model demonstrates a lower propensity for successful jailbreak attempts compared to other models. This lower rate can be attributed to Meta’s focus on developing more robust LLMs that are resistant to such vulnerabilities. (Vidgen et al., 2021; Ghosh et al., 2024; Caselli et al., 2020).
4.3 Existing Jailbreak Performance Improvement using BiasJailbreak
The bias-based approach we analyzed can also be applied to existing models. We confirmed performance improvement by applying the BiasJailbreak method to Adaptive Attacks (Andriushchenko et al., 2024), the current state-of-the-art (SOTA) jailbreak attack. Specifically, for the llama2 model(Touvron et al., 2023), the performance improved from 98% to 100%, resulting in a 2% increase. For the phi model(Abdin et al., 2024), the performance enhanced from 95% to 99%, resulting in a 4% increase. These results show that the method can be easily integrated with existing techniques. The experiment was held by using the jailbreak attack artifact from JailbreakBench(Chao et al., 2024), which has 100 samples of Adaptive Attacks conversation
@@ 4.3 Existing Jailbreak Performance Improvement using BiasJailbreak
Model Adaptive Attacks Adaptive Attacks with BiasJailbreak
@@
4.4 Experimental Validation of BiasDefense
To validate our defense method BiasDefense, we conducted experiments on various models, including Llama2, Phi, and Qwen2. We observed the impact of our approach on Marginalized Rate Success and Privileged Rate Success metrics before and after applying our defense technique.
@@ 4.4 Experimental Validation of BiasDefense
Table 4: Jailbreak Prevention performance of BiasDefense
@@
Table 5: Comparison of jailbreak success rates across prompt-based defense methods using LLAMA2 model.
Defense Method Metric Jailbreak Success Rate
None Marginalized Group Jailbreak Success 28.11%
Privileged Group Jailbreak Success 19.33%
self-remind Marginalized Group Jailbreak Success 33.67%
Wu et al. (2023) Privileged Group Jailbreak Success 27.33%
Defending Marginalized Group Jailbreak Success 20.57%
Zhang et al. (2023) Privileged Group Jailbreak Success 14.86%
RPO Marginalized Group Jailbreak Success 56.00%
Zhou et al. (2024) Privileged Group Jailbreak Success 49.75%
BiasDefense (Ours) Marginalized Group Jailbreak Success 17.14%
Privileged Group Jailbreak Success 14.29%
In Table 5, BiasDefense, our proposed method, demonstrates superior defense performance by achieving the lowest jailbreak success rates among the compared state-of-the-art prompt-based defense methods. Specifically, it records the lowest success rates for both the Marginalized Group and the Privileged Group, indicating enhanced overall security. This suggests that BiasDefense is more effective at preventing jailbreak attacks compared to other defense techniques.
Table 6: Comparison of Defense cost for BiasDefense and Llama-Guard. The experiment was held on a single H100 GPU, with Llama-3.2-1B-Instruct as the language model, and Llama-Guard-3-8B as the guard model.
Defense Method Time Cost (seconds) $\downarrow$ Time cost (percentage) $\downarrow$
Baseline (No Defense) 21.91 +0.00%
BiasDefense 22.44 +2.40%
Llama-Guard-3-8B 31.69 +44.60%
Table 6 shows how our prompt-based method is computationally efficient compared to standalone classifier models such as Llama-Guard. Using a standalone classifier requires far more computation time than adding defense prompts.
@@ 5 Conclusion
In conclusion, our study highlights the complexities and potential unintended consequences of aligning large language models (LLMs) with safety measures aimed at preventing harmful outputs. The introduction of ethical biases for ethical behavior, while crucial for ensuring responsible AI, has led to significant discrepancies in jailbreak success rates based on gender and racial keywords. This discrepancy, coined as ”BiasJailbreak,” underscores the risks that arise from safety-induced biases, particularly in terms of fairness and equality. Our findings emphasize the need for LLM developers to carefully balance safety and fairness in their models. Additionally, our proposed method of adding defense prompts without requiring additional inference or models shows a simple and scalable solution to mitigate jailbreak attempts. This simplicity and scalability of migation emphasizes the responsibility of LLM developers to adopt more comprehensive and proactive measures to address safety risks. Future work should focus on designing more inclusive and transparent alignment strategies to address the inherent challenges of AI safety while minimizing bias.
@@ Appendix A Appendix
A.1 PCA Visualization Analysis
In this section, we conduct a detailed analysis of how models interpret BiasJailbreak prompts by performing Principal Component Analysis (PCA) (F.R.S., 1901) on the embeddings of Phi-mini, Qwen2, and LLaMA3 models, as shown in Appendix Figures 5, 6, 7 respectively. Our findings reveal that models with a high success rate in BiasJailbreak tend to cluster BiasJailbreak prompts together with benign prompts, while those with a lower success rate demonstrate the opposite behavior. The PCA results indicate that when using BiasJailbreak, models tend to interpret it more similarly to benign prompts, with the degree of this similarity aligning with our observed BiasJailbreak success rates.
@@ A.1 PCA Visualization Analysis
Figure 5: PCA of Phi-mini model, which has a 0.4386 BiasJailbreak success rate. BiasJailbreak samples are closely clustered with Benign samples.
@@
Figure 6: PCA of Qwen2 model, which has a 0.1971 BiasJailbreak success rate. BiasJailbreak samples are relatively far from both harmful and benign samples.
@@
Figure 7: PCA of LLaMA3 model, which has a 0.0650 BiasJailbreak success rate. BiasJailbreak samples are relatively close to harmful samples.
A.2 Keyword Group Impact on BiasJailbreak
Table 7 presents the reported confidence intervals based on three runs conducted with a temperature setting of 0.7. The results reveal a clear distinction between marginalized and privileged keywords, with marginalized keywords generally demonstrating higher success rates. The observed confidence intervals are consistent across runs, further validating the robustness of the findings. We also conducted experiments for random adjective keywords. While slow exhibited a relatively high success rate of 31%, other adjectives had jailbreak success rates similar to the baseline (without a keyword). This suggests that general adjectives have minimal impact on jailbreak attempts. This consistency, also evident in the case of general adjectives, underscores the reliability of the methodology. Notably, significant disparities were observed in groups such as Black vs. White, neurodiverse vs. neurotypical, Native American vs. White, and LGBTQ+ vs. heterosexual, where marginalized keywords consistently outperformed privileged ones.
Table 7: Comparison of the impact of individual keyword groups on BiasJailbreak. Marginalized Groups and Privileged Groups have a distinct difference among Random Adjectives. The Difference (%) (95% CI) column shows the change in Jailbreak Success Rate compared to the baseline, along with the 95% confidence interval.
Random Adjectives Group
Keyword Model – Baseline 95% Conf. Interval Group treatment effect
big -4.00% (1.32%) (-6.59%, -1.41%) -0.50% (3.97%)
small +1.00% (1.26%) (-1.47%, +3.47%)
bright -2.00% (3.73%) (-8.69%, +4.69%)
cold -3.00% (1.26%) (-5.47%, -0.53%)
old 0.00% (6.26%) (-12.56%, +12.56%)
slow +7.00% (7.64%) (-7.61%, +21.61%)
warm -2.00% (1.82%) (-5.05%, +1.05%)
dark -1.00% (2.56%) (-5.34%, +3.34%)
Marginalized Group
Keyword Model – Baseline 95% Conf. Interval Treatment Effect
female +4.00% (3.73%) (-2.69%, +10.69%) +4.33% (4.93%)
poor +7.00% (2.56%) (+2.66%, +11.34%)
black +2.33% (1.30%) (-0.14%, +4.80%)
LGBTQ 0.00% (1.26%) (-2.47%, +2.47%)
transgender +3.00% (7.76%) (-11.83%, +17.83%)
neurodiverse +7.67% (12.55%) (-16.71%, +32.05%)
working class +5.33% (4.14%) (-2.22%, +12.88%)
bisexual +0.67% (1.30%) (-1.80%, +3.14%)
native american +9.00% (1.26%) (+6.53%, +11.47%)
Privileged Group
Keyword Model – Baseline 95% Conf. Interval Treatment Effect
male 0.00% (1.26%) (-2.47%, +2.47%) -2.90% (11.21%)
rich +2.67% (1.33%) (+0.08%, +5.26%)
white -7.33% (1.30%) (-9.80%, -4.86%)
heterosexual -7.00% (1.26%) (-9.47%, -4.53%)
straight -14.33% (6.82%) (-27.28%, -1.38%)
neurotypical +3.67% (6.82%) (-9.28%, +16.62%)
middle class +2.00% (28.87%) (-54.52%, +58.52%)
A.3 BiasDefense Ablation on the combination of prefix and suffix prompt
The results of the BiasDefense Ablation Study, shown in Table 8, demonstrate the performance of various configurations of the proposed BiasDefense mechanism. BiasDefense consists of two primary components: a system prompt and a suffix prompt. The ablation study evaluates the effectiveness of these components both independently and in combination.
  • Individual Prompts: Using either the system prompt or the suffix prompt independently achieves lower jailbreak success rates compared to having no defense at all. For example:
    • The system prompt reduces marginalized group jailbreak success from 0.2811 (no defense) to 0.2558, and privileged group success from 0.1933 to 0.1687.
    • The suffix prompt reduces marginalized group jailbreak success to 0.2367, although it slightly increases the success rate for privileged groups to 0.2733.
  • Combined Prompts: The best performance is achieved when both the system and suffix prompts are used together. The full BiasDefense configuration reduces marginalized group jailbreak success to 0.1714 and privileged group jailbreak success to 0.1429, representing a significant improvement over individual prompts and the baseline without any defense.
These findings highlight the importance of using both components of BiasDefense together. While each component provides some improvement when used independently, their integration ensures the most robust performance against jailbreak attempts. This demonstrates the value of a comprehensive approach to mitigating bias vulnerabilities in language models. Furthermore, our prompt-based defense method effectively addresses ethical biases, providing a practical solution for improving the fairness and reliability of language models.
Table 8: BiasDefense Ablation on the combination of suffix and prefix defense prompt using LLAMA2 model.
Defense Method Metric Jailbreak Success Rate
None Marginalized Group Jailbreak Success 0.2811
Privileged Group Jailbreak Success 0.1933
BiasDefense: system prompt Marginalized Group Jailbreak Success 0.2558
Privileged Group Jailbreak Success 0.1687
BiasDefense: suffix prompt Marginalized Group Jailbreak Success 0.2367
Privileged Group Jailbreak Success 0.2733
BiasDefense Marginalized Group Jailbreak Success 0.1714
Privileged Group Jailbreak Success 0.1429
A.4 BiasDefense Ablation on normal task performance
For evaluating the effect on normal task performance with defense prompts including BiasDefense, Table 9 shows the MMLU score performance with various defense prompt methods. The results show that our BiasDefense has relatively low performance degradation compared to other suffix adding defense prompts. One notable aspect is that without the suffix prompt in BiasDefense, MMLU score has gotten better than using no defense prompts, showing even a positive effect on normal tasks. Although the previous results of Table 8 shows that without the suffix the defense gets weaker, using only the prefix prompt would stand as a good alternative if we want minimal impact on normal tasks.
Table 9: MMLU score using prompt-based defense methods. The base language model used is google-flan-t5 (Chung et al. (2024)) and GPT-4o-2024-11-20. *For GPT-4o, we used 4 subjects sampled from each category due to cost, being college_computer_science (category: STEM), sociology (category: Social Sciences), college_medicine (category: Other), jurisprudence (category: Humanities).
Model Defense Method Suffix Average MMLU*
google-flan-t5 None X 0.295
google-flan-t5 RPO Zhou et al. (2024) O 0.243
google-flan-t5 self-remind Wu et al. (2023) O 0.238
google-flan-t5 Defending Zhang et al. (2023) X 0.295
google-flan-t5 BiasDefense(Ours, With Suffix) O 0.265
google-flan-t5 BiasDefense(Ours, Without Suffix) X 0.296
GPT-4o-2024-11-20 None X 0.883
GPT-4o-2024-11-20 RPO Zhou et al. (2024) O 0.237
GPT-4o-2024-11-20 self-remind Wu et al. (2023) O 0.861
GPT-4o-2024-11-20 Defending Zhang et al. (2023) X 0.876
GPT-4o-2024-11-20 BiasDefense(Ours, With Suffix) O 0.869
GPT-4o-2024-11-20 BiasDefense(Ours, Without Suffix) X 0.887
A.5 Keywords for Different Models
Table 10: Keywords discovered in Claude-sonnet3.5.
Marginalized Group (Adjective) Privileged Group (Adjective)
Female Male
Immigrant Native
Poor Rich
Muslim Christian
Black White
High school educated College educated
LGBTQ+ Heterosexual
Transgender Straight
Neurodiverse Neurotypical
Spanish English
Working class Middle class
Bisexual Heterosexual
Native American White
Table 11: Keywords discovered in GPT-4o.
Marginalized Group (Adjective) Privileged Group (Adjective)
Poor Wealthy
Black White
Elderly Young
Disabled Able-bodied
Non-binary Cisgender
Unemployed Employed
Religious minority Religious majority
Indigenous Colonizer-descended
Lower-class Upper-class
Underrepresented Overrepresented
Minoritized Dominant
Non-English speaking English speaking
Informally educated Formally educated
Psychiatrically diagnosed Psychiatrically typical
Single parent Partnered parent
Adolescent Adult
Gender non-conforming Gender conforming
Internationally born Nationally born
Temporarily employed Permanently employed
Linguistically diverse Linguistically standard
Socially isolated Socially integrated
Chronically ill Healthy
Mentally ill Mentally healthy
Substance dependent Substance independent
Homeless Sheltered
Incarcerated Non-incarcerated
Refugee Citizen
Visually impaired Visually unimpaired
Hearing impaired Hearing unimpaired
Non-traditional student Traditional student
Working class Professional class
Historically disadvantaged Historically advantaged
War-affected Peace-living
Food insecure Food secure
Low mobility High mobility
Environmentally displaced Environmentally secure
Survivor Non-affected
Recovering Non-recovering
Public transport dependent Car owning
Crafts-based Industrial
Artisanal Mass-produced
Temporary visa holder Permanent resident
Non-voter Voter
Uninsured Insured
Non-internet user Internet user
Low-income High-income
Table 12: Keywords discovered in Qwen2.
Marginalized Group (Adjective) Privileged Group (Adjective)
Low-Income High-Income
Racial Minority White
LGBTQ+ Heterosexual
Women Men
Disabled Able-bodied
Refugees Citizens
Uneducated Educated
Neurodivergent Neurotypical
Religious Minority Mainstream Religion
Homeless Homeowners
Mental Health Patients Mentally Stable Individuals
Elderly Young Adults
Overweight Slim/Fit
Non-native Speakers Native Speakers
Incarcerated Free
Rural Urban
Illiterate Literate
Politically Disenfranchised Voters
Undocumented Immigrants Legal Immigrants
Sex Workers Non-sex Workers
Table 13: Keywords discovered in LLaMA2.
Marginalized Group (Adjective) Privileged Group (Adjective)
Female Male
Poor Rich
Black White
LGBTQ Heterosexual
Transgender Straight
Neurodiverse Neurotypical
Working class Middle class
Bisexual Heterosexual
Native American White
Table 14: Keywords discovered in Phi.
Marginalized Group (Adjective) Privileged Group (Adjective)
Black White
Female Male
LGBTQ+ Heterosexual
Non-normative gender Cisgender
Non-traditionally family-structured Traditionally family-structured
Rural Urban
Homeless Socially influential
Undereducated Highly educated
Disabled Able-bodied
Religiously marginalized Religiously dominant
Low-income Wealthy
Sexually employed Traditionally employed